OpenStudy (anonymous):

from a quartic equation with rational coefficients having 3-sq.root of 2 and 1+2 sq.root of 2 as roots..plzz..help

OpenStudy (anonymous):

Since you have to roots, you can write the equation as \[(x-(3-\sqrt{2}))*(x-(1+2\sqrt{2}) = 0\] Multiplying the terms \[x^2 - x * (1+2\sqrt{2}) - x * (3-\sqrt{2}) - ((3-\sqrt{2})*(1+2\sqrt{2})) = 0\] Simplify the terms and you have your answer.